# LARA: Linguistic-Adaptive Retrieval-Augmented LLMs for Multi-Turn Intent Classification

### v1

@@ Abstract
Following the significant achievements of large language models (LLMs), researchers have employed in-context learning for text classification tasks. However, these studies focused on monolingual, single-turn classification tasks. In this paper, we introduce LARA (Linguistic-Adaptive Retrieval-Augmented Language Models), designed to enhance accuracy in multi-turn classification tasks across six languages, accommodating numerous intents in chatbot interactions. Multi-turn intent classification is notably challenging due to the complexity and evolving nature of conversational contexts. LARA tackles these issues by combining a fine-tuned smaller model with a retrieval-augmented mechanism, integrated within the architecture of LLMs. This integration allows LARA to dynamically utilize past dialogues and relevant intents, thereby improving the understanding of the context. Furthermore, our adaptive retrieval techniques bolster the cross-lingual capabilities of LLMs without extensive retraining and fine-tune. Comprehensive experiments demonstrate that LARA achieves state-of-the-art performance on multi-turn intent classification tasks, enhancing the average accuracy by 3.67% compared to existing methods.
Keywords:
In-Context Learning LLM Multi-turn text classification
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Chatbots are an essential tool that automatically interacts or converses with customers. It plays a crucial role for international e-commerce platforms due to the rising consumer demand for instant and efficient customer service. Chatbots represent a critical component of dialogue systems [24] that can answer multiple queries simultaneously by classifying intent from the user’s utterance to reduce waiting times and operational costs. Naturally, the interaction with users could turn into a multi-turn conversation if they require more detailed information about the query. Developing an intent classification model for a dialogue system isn’t trivial, even if it’s a typical text classification task. As we must consider contextual factors such as historical utterances and intents, failing to understand the session context while recognizing the user intention usually leads to more visible errors, as it would invoke a completely wrong application or provide an unrelated answer [28]. It’s not surprising that it faces several challenges in dialogue understanding.
@@
Figure 1: Annotation challenge of multi-turn intent classification dataset
@@
The biggest challenge is that the multi-turn dataset is hard to collect. Some studies have already been done on this problem of dialogue understanding in multi-turn intent classification [18, 25, 16]. However, they are made under the assumption of the availability of multi-turn training data, which is usually not the case in the real world.
@@
Unlike emotion recognition in conversation (ERC) with only less than 10 classes or topic classification within dialogue state tracking (DST) with tens of topics, there are hundreds of intents within the knowledge base of a chatbot to cover users’ specific intents in each market, it increases the complexity of classification tasks and multi-turn data annotation. Annotators can easily make mistakes and spend more time making decisions due to the numerous intents. Combined, these make it a high-cost and time-consuming annotation task, and it’s unrealistic to annotate large-scale multi-turn datasets manually. However, the performance will most likely suffer without enough training sample size. This calls for a more efficient method in solving the challenge [13].
@@
To tackle the above challenge, we propose Linguistic-Adaptive Retrieval-Augmentation, or LARA, which offers a pipeline of techniques to adopt only single-turn training data to optimize multi-turn dialogue classification. LARA first leverages an XLM-based model trained on single-turn classification datasets for each market, thus simplifying data construction and maintenance. Subsequently, LARA advances the field by selecting plausible candidate intents from user utterances and employing a retriever to gather relevant questions for prompt construction. This process facilitates in-context learning (ICL) with multi-lingual LLMs (MLLMs), significantly enhancing model efficacy without the need for market-specific multi-turn models.
@@
    We introduce LARA to effectively address multi-turn data collection issue through XLM-based model training and ICL with MLLM.
@@
    We conduct experiments on our e-commerce multi-turn dataset across six languages, showing that LARA model achieves state-of-the-art results and reduces inference time during ICL with MLLMs.
@@ 2 Related Work
2.1 Intent Classification
Intent classification is a typical text classification task where class labels are intent names. Various neural networks have inspired a wide range of studies aimed at creating neural models for text classification. Various neural model structures such as CNN [5, 8], LSTM [29] and GCN [30, 11] have proven more effective than conventional methods [9, 23] based on statistical features. Additionally, some studies [34, 35] utilize label embeddings and train them simultaneously with the input texts. In recent developments, the accomplishments of large-scale pre-training language models (PLM) [7] have generated significant interest in incorporating this pre-training approach [17] into monolingual and multilingual text classification [4]. This has resulted in substantial advancements in few-shot [3] and zero-shot learning [31]. However, most works are single-turn text classification tasks, which is unsuitable for multilingual multi-turn intent classification.
2.2 Modeling Multi-turn Dialogue Context
Modelling the multi-turn dialogues is the foundation for dialogue understanding tasks. Previous works adopt bidirectional contextual LSTM [6] to create contextual-aware utterance representation on MultiWOZ intent classification [2]. Recent works use PLM as a sentence encoder [20] on emotion recognition in conversation(ERC). Specifically, [10] used PLM to encode the context and speaker’s memory and [15] enhance PLM by integrating multi-turn info from the utterance, context and dialogue structure through fine-tuning. However, all of their tasks adopt the multi-turn dialogue training set, which is hard to collect for an e-commerce chatbot. Our method attempts to combine an XLM-based model trained on the single-turn dataset into an in-context retrieve augmented pipeline with LLM, solving the multi-turn intent classification task in a zero-shot setting.
2.3 In-context Retrieval
In-context learning (ICL) with LLM like GPT-3 [1] demonstrates the significant improvement on few-shot/zero-shot NLP tasks. ICL has been successful in tasks like semantic parser [14, 21], intent classification [32] and other utterance-level tasks. Some researchers [27, 12] also apply ICL for dialogue state tracking, but it didn’t perform as well as other methods. Future studies might look into better ways to retrieve dialogues and improve tasks’ setup. Retrieval part, Most research on in-context learning (ICL) usually deals with single sentences or whole documents, but we are interested in finding and understanding dialogues. Generally, there are two types of systems to find the relevant dialogues: the first is LM-score based retrieval. They [19, 22] check the probability of a language model, like GPT-3, is to decode the right answer based on an example. The second type defines similarity metrics between task results and uses them as the training objective for the retriever. Both K-highest and lowest examples are used as positive and negative samples to help the system learn. The most pertinent research on dialogue retrieval concentrates on areas such as knowledge identification [26] and response selection [33]. Our objectives and settings differ from them.
@@ 3 Problem Formulation
3.1 Single-turn Intent Classification
In the single-turn scenario, the objective is to classify a user’s query $q$ into one of the predefined intent classes =$\mathcal{I}=\{I_{i}\}_{i=1}^{k}$, where $k$ often exceeds 200. Queries can range from informational requests to action-oriented dialogues. Recognizing the correct intent enables the dialogue system to provide relevant solutions or perform specific actions. This process is further complicated in transnational applications that must accommodate queries in multiple languages and adapt to localized business operations.
@@ 3.2 Multi-turn Intent Classification
Multi-turn scenarios involve a series of user queries $\mathcal{Q}=\{q_{i}\}_{i=1}^{n}$, with the aim of identifying the intent of the final query $q_{n}$. Unlike single-turn recognition, multi-turn recognition must account for the entire conversational context $\mathcal{C}$, which includes historical queries and their corresponding intents. This context-dependency introduces additional complexity, requiring models to interpret nuanced conversational dynamics and adjust to evolving user intentions over the course of an interaction.
@@ 3.3 Objective
This work aims to devise a methodology that leverages the readily available single-turn training data to address the inherent challenges of multi-turn intent recognition without the need for extensive multi-turn dataset curation. By proposing a model that dynamically adapts to the conversation’s context while minimizing reliance on large-scale annotated multi-turn datasets, we seek to mitigate the significant annotation challenges and enable more effective intent recognition in complex multi-turn dialogues.
@@ 4 LARA: Linguistic-Adaptive Retrieval-Augmentation
The LARA framework addresses the multi-turn intent recognition challenge through zero-shot in-context learning with single-turn demonstrations, guided by a crafted instruction prompt. First, a single-turn intent classification model $\mathcal{M}_{c}$ is used to narrow down the intents to be included in the ICL prompt, which hereby referred to as candidate intents. This step is necessary due to the limited LLM context window, and it also helps to filter out extra noises from direct demonstration retrieval. Then, for every candidate intent, in-context demonstrations are selected by retrieving single-turn examples that are semantically-similar to the multi-turn test sample. Finally, an instruction prompt for multi-turn intent recognition is formulated by combining the demonstrations and test user queries.
Algorithm 1 Candidate Intent Selection
$\mathcal{Q}_{c}=\{q_{n},q^{1}_{n},...,q^{n-1}_{n}\}$
for  each item $q_{i}$ in $\mathcal{Q}_{c}$  do
     /* Get embedding H of [CLS] token */
     $H_{i}=\Phi_{XLMR}(q)^{[CLS]}\in\mathbb{R}^{d}$, $d$ is the hidden dimension
     /* Pass the embedding through a linear layer to get class probability P */
     $P_{i}=softmax(H\cdot W_{c}+b_{c})$, $W_{c}\in\mathbb{R}^{d\times|\mathcal{I}|}$ and $b_{c}\in\mathbb{R}^{|\mathcal{I}|}$
     /* Select the intent with highest probability */
     $I_{i}=argmax(P_{i})$
end for
return $\mathcal{I}_{c}=\{I_{n},I^{1}_{n},...,I^{n-1}_{n}\}$
@@
     $X_{i}=get\_training\_samples\_from\_\mathcal{D}\_for\_intent(I_{i})$
     /* get embedding for each training sample */
@@
     $S_{i}=\{cosine\_similarity(H_{q},h_{j})\}_{j=1}^{|H_{X_{i}}|},h\in H_{X_{i}}$
     /* select nearest demonstrations */
@@
4.0.1 Single-turn Intent Recognition Model ($\mathcal{M}_{c}$)
A text classification model is trained on the annotated single-turn dataset $\mathcal{D}$. Given a query $q$, we adopt the [CLS] token embedding from XLM-RoBERTa-base model with weight $\Phi_{XLMR}$ as the text representation $H$. $\Phi_{XLMR}$ had been further pretrained with contrastive learning to give meaningful representation for [CLS] token. $H$ is then fed into a linear layer along with a softmax function to obtain intent probabilities. The intent with the highest probability is used.
During candidate intents selection, the last query $q_{n}$ is first combined with each historical query in $\mathcal{C}$ to form a query combination set $\mathcal{Q}_{c}$ = $\{q_{n},q^{1}_{n},...,q^{n-1}_{n}\}$, where $q^{i}_{n}$ means the text concatenation of $q_{i}$ with $q_{n}$ using a comma. $\mathcal{M}_{c}$’s inference on these combinations yields the candidate intents $\mathcal{I}_{c}$. The selection process is detailed in Algorithm 1.
4.0.2 Retrieval Augmentation
$\Phi_{XLMR}$ is also utilized here to gather demonstrations for each intent in $\mathcal{I}_{c}$ based on their cosine similarity to test queries. Here, the demonstrations refer to a sequence of annotated examples that provide LLM with decision-making evidence and specify an output format for natural language conversion into labels during ICL. The details of ICL demonstration retrieval is included in Algorithm 1.
4.0.3 Prompt Construction and LLM Inference
The task instruction $\mathcal{T}$, combined with demonstrations $\mathcal{E}$, conversational context $\mathcal{C}$, and the query $q_{n}$, forms the input prompt $\mathcal{P}$ for the LLM. To accommodate real-time application latency requirements, two additional methods were explored to constrain the model to generate single-token symbols representing intents, detailed as $\mathcal{P}_{symbolic}$ and $\mathcal{P}_{prepend}$, with examples provided in the appendix. Model outputs are greedily decoded, ensuring efficient and accurate intent recognition.

### v2

# LARA: Linguistic-Adaptive Retrieval-Augmentation for Multi-Turn Intent Classification

@@ Abstract
Multi-turn intent classification is notably challenging due to the complexity and evolving nature of conversational contexts. This paper introduces LARA, a Linguistic-Adaptive Retrieval-Augmentation framework to enhance accuracy in multi-turn classification tasks across six languages, accommodating a large number of intents in chatbot interactions. LARA combines a fine-tuned smaller model with a retrieval-augmented mechanism, integrated within the architecture of LLMs. The integration allows LARA to dynamically utilize past dialogues and relevant intents, thereby improving the understanding of the context. Furthermore, our adaptive retrieval techniques bolster the cross-lingual capabilities of LLMs without extensive retraining and fine-tuning. Comprehensive experiments demonstrate that LARA achieves state-of-the-art performance on multi-turn intent classification tasks, enhancing the average accuracy by 3.67% from state-of-the-art single-turn intent classifiers.
LARA: Linguistic-Adaptive Retrieval-Augmentation for Multi-Turn Intent Classification
Junhua Liu1,3,, Yong Keat Tan2,, Bin Fu2,, Kwan Hui Lim3 1Forth AI 2Shopee 3Singapore University of Technology and Design
**footnotetext: Equal Contributions.$\dagger$\$$\dagger$\$footnotetext: Correspondence: bin.fu@shopee.com
@@ 1 Introduction
A chatbot is an essential tool that automatically interacts or converses with customers. It plays a crucial role for international e-commerce platforms due to the rising consumer demand for instant and efficient customer service. Chatbots represent a critical component of dialogue systems Weld et al. (2021) that can answer multiple queries simultaneously by classifying intent from the user’s utterance to reduce waiting times and operational costs. Naturally, the interaction with users could turn into a multi-turn conversation if they require more detailed information about the query. Developing an intent classification model for a dialogue system is not trivial, even if it is a typical text classification task. As we must consider contextual factors such as historical utterances and intents, failing to understand session context while recognizing user intention often leads to visible errors. It would invoke a completely wrong application or provide an unrelated answer Xu and Sarikaya (2014). As a result, it faces several challenges in dialogue understanding.
@@
Figure 1: Annotation pipeline of multi-turn intent classification dataset
@@
The biggest challenge is that multi-turn datasets are hard to collect. While there are some studies on dialogue understanding in multi-turn intent classification Ren and Xue (2020); Wu et al. (2021a); Qu et al. (2019), they are made under the assumption of the availability of multi-turn training data, which is usually not the case in the real world.
@@
Figure 1 shows the annotation pipeline of multi-turn intent classification. Unlike emotion recognition in conversation (ERC) with only less than 10 classes or topic classification within dialogue state tracking (DST) with tens of topics, there are hundreds of intents within the knowledge base of a chatbot to cover users’ specific intents in each market, which increases the complexity of classification tasks and multi-turn data annotation. Annotators can easily make mistakes and spend more time making decisions due to the numerous intents. Combined, these make it a high-cost and time-consuming annotation task, and it is unrealistic to annotate large-scale multi-turn datasets manually. However, the performance will most likely suffer without enough training sample size. This calls for a more efficient method in solving the challenge Mo et al. (2023).
@@
To tackle the above challenge, we propose Linguistic-Adaptive Retrieval-Augmentation, or LARA, which offers a pipeline of techniques to adopt only single-turn training data to optimize multi-turn dialogue classification. LARA first leverages an XLM-based model trained on single-turn classification datasets for each market, thus simplifying data construction and maintenance. Subsequently, LARA advances the field by selecting plausible candidate intents from user utterances and employing a retriever to gather relevant questions for prompt construction. This process facilitates in-context learning (ICL) with multi-lingual LLMs (MLLMs), significantly enhancing model efficacy without requiring market-specific multi-turn models.
@@
    We introduce LARA as a multi-turn classification model only leveraging single-turn datasets to effectively address multi-turn data collection issues.
@@
    We conduct experiments on our e-commerce multi-turn dataset across six languages. LARA achieves state-of-the-art results and reduces inference time during ICL with MLLMs.
@@ 2 Related Work
Modeling Multi-turn Dialogue Context: Modelling the multi-turn dialogues is the foundation for dialogue understanding tasks. Previous works adopt bidirectional contextual LSTM Ghosal et al. (2021) to create context-aware utterance representation on MultiWOZ intent classification Budzianowski et al. (2018). Recent works use PLM as a sentence encoder Shen et al. (2021) on emotion recognition in conversation (ERC). Specifically, Lee and Lee (2022) used PLM to encode the context and speaker’s memory and Qin et al. (2023) enhance PLM by integrating multi-turn info from the utterance, context and dialogue structure through fine-tuning. However, all of their tasks adopt the multi-turn dialogue training set, which is hard to collect for an e-commerce chatbot. Our method attempts to combine an XLM-based model trained on the single-turn dataset into an in-context retrieval augmented pipeline with LLM, solving the multi-turn intent classification task in a zero-shot setting.
In-context Retrieval: In-context learning (ICL) with LLM like GPT-3 Brown et al. (2020) demonstrates the significant improvement on few-shot/zero-shot NLP tasks. ICL has been successful in utterance-level tasks like intent classification Yu et al. (2021). As for the Retrieval part, most research on in-context learning (ICL) usually deals with single sentences or document retrieval, but we are interested in finding and understanding dialogues. Generally, there are two types of systems to find the relevant dialogues: the first is LM-score based retrieval. They Rubin et al. (2021); Shin et al. (2021) check the probability of a language model, like GPT-3, to decode the right answer based on an example. The second type defines similarity metrics between task results and uses them as the training objective for the retriever. Both K-highest and lowest examples are used as positive and negative samples to help the system learn. The most pertinent research on dialogue retrieval concentrates on areas such as knowledge identification Wu et al. (2021b) and response selection Yuan et al. (2019). Our objectives and settings differ from them.
@@ 3 Problem Formulation
3.1 Hierarchical Text Classification
Hierarchical text classification (HTC) is a type of text classification where the classes or categories are organized in a hierarchy or tree structure. Instead of having a flat list of categories, the categories are arranged in a nested manner, so we need to consider the relationships of the nodes from different levels in the class taxonomy.
The intents in our scenario are organised in a hierarchical tree structure. Specifically, each category can belong to at most one parent category and can have arbitrary number of children categories. Our class taxonomy $\mathcal{T}$ is a tree with fixed depth 3 and one meta root node, that is, the distance from the root node to all leaf nodes is 3, and the unique path from the root node to each leaf node will form one intent.
We formulate the single-turn HTC task in our scenario as such, given a text $q_{i}$, the purpose is to predict a subset $I$ of the complete intent set $\mathcal{I}$, where the size of the subset $|I|$ containing the category from each layer is 3 (excluding root node). Generally, the number of intent sets exceeds 200.
@@ 3.2 Multi-turn Intent Classification
Multi-turn scenario shares the same $\mathcal{T}$ and $\mathcal{I}$ as single-turn scenario. It involves a series of user queries $\mathcal{Q}=\{q_{i}\}_{i=1}^{n}$ in dialogues, the objective is to identify the intent of the final query $q_{n}$. Multi-turn recognition must account for the entire conversational context $\mathcal{C}=\{q_{i}\}_{i=1}^{n-1}$, which includes the historical queries. This context-dependency introduces additional complexity, requiring models to interpret nuanced conversational dynamics and adjust to evolving user intentions over the course of an interaction.
@@ 3.3 Objective
This work aims to use easily accessible single-turn data to improve the accuracy of multi-turn intent recognition without requiring any multi-turn training datasets.
@@ 4 LARA: Linguistic-Adaptive Retrieval-Augmentation
The LARA framework shown in Figure 2 addresses the multi-turn intent recognition challenge through zero-shot in-context learning with single-turn demonstrations guided by a crafted instruction prompt. First, a single-turn classification model $\mathcal{M}_{c}$ is trained on single-turn dataset and used to narrow down the intents to be included in the ICL prompt, which are henceforth referred to as candidate intents. This step is necessary due to the limited LLM context window, and it also helps to filter out extra noise from direct demonstration retrieval. Then, for every candidate intent, in-context demonstrations are selected by retrieving single-turn examples that are semantically similar to the multi-turn test sample. Finally, an instruction prompt for multi-turn classification is formulated by combining the demonstrations and test user queries.
4.1 Single-turn Classification Model ($\mathcal{M}_{c}$)
Before diving into LARA, we must train a single-turn hierarchical text classification model on our single-turn dataset $\mathcal{D}$. The model is an ensemble of a simple label-attention model Zhang et al. (2020) and a state-of-the-art HTC approach named HiTIN Zhu et al. (2023). The label-attention model only considers the semantic info of user utterances and label-query attention info, which ignores the hierarchical tree-like structure in our intent system. So, we build model $\mathcal{M}_{c}$, integrating taxonomic structure via the tree isomorphism network within HiTIN into our label-attention model. More details are provided in Appendix A.
We conduct some experiments on our multilingual dataset. The result in Table 1 shows that HiTIN alone performs better than the label-attention model, and ensembling the two methods further improves performance on average scores. We will use this model as our baseline and generate candidate intents within LARA.
BR ID MY PH SG TH TW VN avg
Traffic weight 150k 212k 27k 46k 5k 36k 36k 43k
Label-attention model 81.55% 87.91% 83.73% 73.29% 83.69% 83.11% 71.55% 74.44% 82.30%
HiTIN 81.85% 89.41% 84.57% 74.75% 84.17% 83.67% 75.38% 75.99% 83.53%
Ensembled 82.66% 89.62% 85.33% 75.98% 84.57% 83.89% 75.19% 75.52% 83.94%
Table 1: Noticeable improvement from adding a state-of-the-art HTC approach.
4.2 Candidate Intents Selection
Query Combination: After receiving the context from user side, the last query $q_{n}$ is first combined with each historical query in conversational context $\mathcal{C}$ to form a query combination set $\mathcal{Q}_{c}$ = $\{q_{n},q^{1}_{n},...,q^{n-1}_{n}\}$, where $q^{i}_{n}$ means the text concatenation of $q_{i}$ with $q_{n}$ using a comma.
Candidate Intent Recognition: $\mathcal{M}_{c}$ predicts few candidate intents $\mathcal{I}_{c}$ from all intents $\mathcal{I}$ on these query combinations $\mathcal{Q}_{c}$. The $\mathcal{M}_{c}$ is then used to perform inference on each of the concatenated queries to get a set of candidate intents $\mathcal{I}_{c}=\{I_{q_{n}},I_{q^{1}_{n}},...,I_{q^{n-1}_{n}}\}$. Finally, we will select the top 3 intents with the highest scores. Note that $\mathcal{I}_{c}$ is a set, so duplicated intents will be removed, and the maximum size of $\mathcal{I}_{c}$ is equal to the number of queries $n$ in a session.
4.3 ICL Retrieval
However, not all training examples under the candidate’s intent $\mathcal{I}_{c}$ will be used in the prompt as demonstrations. Here, the demonstrations refer to a sequence of annotated examples that provide LLM with decision-making evidence and specify an output format for natural language conversion into labels during ICL. Our strategy is to sample training examples that are similar to the test sequence for each candidate intent, a method introduced by Liu et al. (2021). In this process, we first concatenate all the queries in the session with a comma to get the test sequence $q_{all}$. Then, $q_{all}$ is mapped to a vector $H_{q}$ using the [CLS] token embedding from a pre-trained sentence encoder, $\Phi_{XLMR}$. After that, we retrieve all the single-turn training samples for each candidate intent from $\mathcal{D}$ using a sample retriever and also encode them using $\Phi_{XLMR}$. With $H_{q}$ as query, we use an embedding retriever to search through the sampled examples to retrieve top $K$ 1 nearest data examples of each candidate intent based on their cosine similarity to $H_{q}$. Together with the representative query of each candidate intent, the retrieved single-turn samples of all candidate intents form the demonstrations $\mathcal{E}$ for in-context learning. We show the detailed algorithm in 1 below.
4.4 Prompt Construction and LLM Inference
A task instruction $T$ is hand-crafted to guide the model to perform multi-turn intent recognition task by referring to the single-turn demonstrations. The task instruction $T$, combined with demonstrations $\mathcal{E}$, conversational context $\mathcal{C}$, and the query $q_{n}$, forms the input prompt $\mathcal{P}$ for the LLM. The concatenation of each prompt component into one long text is shown in the appendix C. To accommodate real-time application latency requirements, two additional methods were explored to constrain the model to generate single-token symbols representing intents, detailed as $\mathcal{P}_{symbolic}$ and $\mathcal{P}_{prepend}$, with examples also provided in the appendix. Model outputs are greedily decoded, ensuring efficient and accurate intent recognition.
@@ 4.4 Prompt Construction and LLM Inference
     /* Sample retriever: Get the single-turn training samples under intent $I_{i}$/
     $X_{i}=sample\_from\_\mathcal{D}\_for\_intent(I_{i})$
     /* Embedding retriever: Get embedding for each training sample */
@@
     $S_{i}=\{cosine\_similarity(H_{q},h_{j})\}_{j=1}^{|H_{X_{i}}|}$, where $h\in H_{X_{i}}$
     /* Get top $K$ nearest demonstrations */